calculate the osmotic pressure of 1% of kcl solution which dissociate to an extent of 80%

Grade:12

2 Answers

Arun
6 years ago
Dear Vikash
 
KCl ------>  K +  Cl
1 - @.        @ .    @
 
i = 1 - @ +@ +@ = 1 + @ = 1 + 0.8 = 1.8
 
C = 1000/( 74.5* 100)
 = 0.34
 
T = 273 + 27 = 300 K (there is no temp given in question I have assumed temp as 27°c)
 
Now
 
\pi = i C RT = 1.8 * 0.34 * 0.0821 * 300
   = 5.98 atm
 
 
Regards
Arun (askIITians forum expert)
 
Adarsh
6 years ago
Dear vikash,
osmotic pressure = i.C.R.T where i= vant hoff factor , c is conc. , R is gas constant and T is temp.
now in your question temperature is not given and data is insufficient to find the concentration , if you have all data then you can calculate osmotic pressure by using above expression..
Thanks and Regards 
Adarsh